INAF queue observing, 2018 December 13-14

Observer: R. Carini, A. Rossi <br />
Telescope Operator: S. Allanson <br />
Support Astronomer: B. Rothberg <br />
Instrument: LBC

Summary

Night duration (12\xB0 el twilight) 1:15-13:15 UT = 12h 00m

Observing time: 11h 40min (97%)

Weather time loss: 0h m (0%)

technical time loss 0h 20m (3%)

Data Summary

At the sunset the sky is clear.

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Nisini telluric Field 3 LUCI1,2 K   luci1.20181214.0069-0072==; luci2.20181214.0016-0019; luci1.20181214.0073-0076; luci2.20181214.0020-0023 lambda=2.22 micron and lambda=2.20 micron respectively
Moretti A6 PSO315 MODS 1,2 red grating 20m/60m mods1r.20181214.0006; mods2r.20181214.0006  
Mancini HATP 48 MODS 1,2 dual grating 6h/5.7h mods1r.0010-0141,mods1b.0002-0118; mods2r.0010-0238,mods2b.0002-0138 completed
Gilli mask 44 MODS1,2 dual grating (4+x)h/4h mods1r.0145-0154,mods1b0.0119-0125; mods2r.0242-0251,mods2b.0139-0145 completed, images mods1r.0147, mods1b.0121, mods1r.0244, mods2b.0142 are not completed
Gilli mask 31 MODS1,2 dual grating 1.33h/4h mods1r.0158-0159, mods1b-0126-0127; mods2r.0255-0256, mods2b.0146-0147  
STD Feige 34 MODS1,2 dual grating     Gilli
CALIBRATION BIAS MODS1,2 2x2 bin     Mancini

Details

*during the sunset the TO needs to run AO checkout

01:06 UT Slew to Telluric for Nisini field 3
  • acquisition luci1.20181214.0062-0068; luci2.20181214.0008-0015
  • during the acquisition we did the mask alignment and the connection lost also.
  • seeing 1.2 on guider
  • sci luci1.20181214.0069-0072; luci2.20181214.0016-0019 lambda=2.22 micron
  • sci luci1.20181214.0073-0076; luci2.20181214.0020-0023 lambda=2.20 micron
  • 01:54 seeing on the dimm 1.5"- 1.6"

01:59 the seeing is to variable to do AO, we switch to MODS

02:23 DIMM seeing 1.2"; we start with a target A6 of Moretti's program
  • acq mods1r.20181214.0002-0005; mods2r.20181214.0002-0005
  • sci mods1r.20181214.0006; mods2r.20181214.0006
  • 02:29 DIMM seeing 1.1"-1.2"
  • 02:45 DIMM seeing 1.0"-1.1"

03:02 slew to the target HATP48, Mancinni's program
  • acq mods1r.20181214.0007-0009; mods2r.20181214.0007-0009
  • 03:16 start the science mods1r.0010-0141,mods1b.0002-0118; mods2r.0010-0238,mods2b.0002-0138
  • 03:27 dimm seeing 1.0"-1.1"
  • MODS1b eposure hags upon completions between 6:15 ut and 6:58 ut (bewteen images mods1b.20181214.0069.fits and mods1b.20181214.0070.fits )
  • dimm 0.9"
  • during the integration the the dimm seeing has been very good 0.8"-0.9"
  • 09:05 stop the integration

09:05 slew to Gilli's program, mask 44
  • acq mods1r.00142-0144; mods2r.0239-0241 dimm seeing 0.8"
  • sci mods1r.0045-0147,mods1b0.0119-0121; mods2r.0242-0245,mods2b.0139-0142
  • 09:25 dimm seeing 0.7"
  • 09:56 lost the connection, dimm seeing about 0.7"
  • 09:58 connection is back
  • in MODS1b there isn't the spectrum of the slit above the alignment holes.the problem is that when you put slits outside the region normally used, you get problems like this. In mods2b we there is the spectrum because MODS 1 and MODS2 are not exactly the same. there is a tilt when spectra are dispersed onto the detector and the tilit is slightly different between blue and red.
  • 10:17 lost the guide in MODS1, we stop the integration, we save the image mods1r.0147 and mods1b.0121 after about 500 sec from the beginnin of the integration.
  • 10:19 we have to acquire again in mods1, but there are problems in the preset
  • 10:20 now the preset works
  • acq mods1r.0148-0149
  • 10:30 again problem wiht tha copointing limit (during modsAlign)
  • we try again the modsAlign
  • acq mods1r.0150
  • 10:35 start again the integration, sci mods1r.0151-0154,mods1b-0122-0125-
  • 10:45 mods2 doesn't guide, we stop the integration and save the images mods1r.0244, mods2b.0142
  • 10:47 acquisition again wiht mods2
  • acq mods2r.0246-0248
  • 11:00 science again mods2r.0249-0251,mods2b.0143-0145
  • 11:04 dimm seeing 0.7"-0.8"
  • 11:46 dimm seeing 0.8"

12:09 we slew to the maks 31 (Gilli)
  • acq mods1r.0155-0157, mods2r.0252-0254
  • 21:23 sci mods1r.0158-0159, mods1b-0126-0127; mods2r.0255-0256, mods2b.0146-0147
  • dimm seeing 0.9"

13:11 slew to std Feige 34 dual, redgrating and red prism (for Gilli and Moretti and De Carli)
  • acq mods1r.0160-0162, mods2r.0257-0260
  • 13:18 send preset again
  • 13:27 send sci
  • dualgrating mods1r.0163-0165, mods1b-0128-0130; mods2r.0261-0263, mods2b.0148-0150
  • redgrating mods1r.0166-0167, mods2r.0264-0266 no good sky too bright

MODS1,2 CALIBRATIONS
  • BIAS 2x2 bin per Mancini
Topic revision: r10 - 15 Dec 2018, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback